OpenStudy (anonymous):

How do I make (1,4) and slope=-2 into a slope intercept form

OpenStudy (wilsondanielle):

Plug the points into the equation and solve for the intercept. Eg: Point (2,6) and slope = -1 y=mx+b 6=-1*2+b 6=-2+b 6+2=b 8=b therefore the slope intercept form would be y=-x+8